OpenStudy (anonymous):

find price (p) that maximizes revenue q(p)=-p^2+24p

OpenStudy (paxpolaris):

using Calculus ???

OpenStudy (anonymous):

yes

OpenStudy (callisto):

q'(p) = -2p + 24 Put q'(p) = 0 -2p + 24 = 0 p = 12 Max. revenue = q(12) = -(12)^2+24(12) = ...